# AQORA: A Fast Learned Adaptive Query Optimizer with Stage-Level Feedback for Spark SQL

### v2

@@ VII-D Ablation Study
In this ablation study, we will try to replace AQORA’s core components with other possible substitutes to test out the effectiveness of AQORA’s design on ExtJOB. AQORA uses PPO as the default reinforcement learning algorithm, a TreeCNN architecture for both actor and critic and action networks, and an action space that includes CBO actions, leading actions (with only one allowed leaf node), and no-op actions. Learning is conducted through the curriculum process described in Sec.V-B3.
VII-D1 Different Reinforcement Learning Algorithms
@@
To evaluate the impact of different reinforcement learning algorithms on query optimization performance, we compare the PPO algorithm used in AQORA with DQN [39]. As shown in Fig. 11(a), AQORA with PPO surpasses Spark SQL in execution latency significantly earlier during training compared to its DQN-based variant. Also, after training on 2400 queries, the PPO-based model achieves consistently lower execution latency than the DQN-based model. Although DQN benefits from experience replay, it suffers from limitations in environments with large action spaces and non-stationary dynamics.
@@
Figure 11: Query performance on ExtJOB with different substituted modules, under varying numbers of training queries.
VII-D2 Different Neural Network Structures
@@
We tested multiple neural network architectures for the actor and critic networks, with the basic model statistics and optimization overhead summarized in Tab. III. Among them, FCNN is the only model that fails to outperform Spark SQL even after training on 2400 queries. Specifically, LSTM exhibits the lowest optimization overhead (251 ms), making it more efficient during both training and inference. However, TreeCNN demonstrates better convergence behavior and achieves higher overall performance after training, albeit with a slightly increased overhead.
TABLE III: Parameters and optimization overhead of different model structure on JOB and ExtJOB dataset
Model Parameters Optimization overhead
TreeCNN 147506 317 ms per query
LSTM 4800 251 ms per query
QueryFormer 879602 671 ms per query
FCNN 410546 432 ms per query
VII-D3 Different Learning Strategy
@@
To evaluate the effectiveness of our learning strategy design, we conducted an ablation study using two alternative strategies (Fig. 11(c)). In the first strategy, we removed the step limit during optimization while retaining the curriculum learning schedule. In the second strategy, we limited the maximum number of steps but removed the curriculum constraint, allowing AQORA to explore the full action space from the beginning.
@@
Experimental results show that the first strategy performs comparably to the default configuration during the early training phase. However, as the curriculum progresses to its second stage, the model tends to take significantly more steps, making it difficult to repeatedly make correct decisions. The second strategy, despite enforcing a strict step limit, ultimately fails to outperform Spark SQL. This is because, in most cases, only a few targeted corrections are sufficient to reach an efficient plan. Without early-stage guidance, the model explores a large action space too freely, which hinders convergence.
VII-D4 Different Action Space
@@
AQORA’s planner extension supports a wide range of action types, we conducted ablation studies to identify which subsets of the action space contribute most to learning efficiency and convergence. As shown in the results, enabling the broadcast hint often leads the model to generate plans that broadcast excessively large tables during training, resulting in significant performance fluctuations and instability. When leading actions are disabled, AQORA loses its ability to explicitly manipulate join ordering, thereby reducing its optimization power and leading to suboptimal plans for many complex queries. Finally, disabling the CBO action prevents AQORA from jumping to an entirely different planning path starting from a CBO-generated plan. This restriction increases the difficulty for AQORA to escape local optima and discover globally better execution plans.
VIII Conclusion
We propose AQORA, the first reinforcement learning-based optimizer on Spark SQL that supports execution-time adaptation through four key innovations: (1) realistic feature encoding, (2) stage-level feedback with real-time intervention, (3) automatic strategy adaptation, and (4) seamless integration via Spark’s extension interface. By dynamically constructing bushy joins through node swapping and broadcast tuning, AQORA reduces Spark SQL’s execution latency by up to 70%, while achieving up to 99.8% faster optimization compared to enumeration-based learning query optimizers.
IX AI-Generated Content Acknowledgement
Portions of this paper’s text were refined using ChatGPT 5 to improve the clarity, grammar, and conciseness of the language. The AI system was used only for linguistic polishing and did not contribute to the conceptual development, technical design, experimental results presented in this work.

### v3

@@ 6.5. Ablation Study
In this ablation study, we will try to replace LQRS’s core components with other possible substitutes to test out the effectiveness of LQRS’s design on ExtJOB. We choose ExtJOB because it is relatively small, making experiments more convenient to run, yet it is still sufficient to reveal performance trends. LQRS uses PPO as the default reinforcement learning algorithm, a TreeCNN architecture for both actor and critic and action networks, and an action space that includes init action, lead action (with only one allowed leaf node), and no-op action. Learning is conducted through the curriculum process described in Section 4.2.2.
6.5.1. Different Reinforcement Learning Algorithms
@@
To evaluate the impact of different reinforcement learning algorithms on query optimization performance, we compare the PPO algorithm used in LQRS with DQN  (Mnih et al., 2015). As shown in Figure 9(a), LQRS with PPO surpasses Spark SQL in execution latency significantly earlier during training compared to its DQN-based variant. Also, after training on 2400 queries, the PPO-based model achieves consistently lower execution latency than the DQN-based model. Although DQN benefits from experience replay, it suffers from limitations in environments with large action spaces and non-stationary dynamics.
@@
Figure 9. Query performance on ExtJOB with different substituted modules, under varying numbers of training queries.
6.5.2. Different Neural Network Structures
@@
We experimented with several model architectures proposed in prior work, including a fully connected neural network (FCNN), TreeCNN  (Marcus et al., 2021), LSTM  (Zhu et al., 2015), and QueryFormer  (Zhao et al., 2022). To adapt these models for LQRS, we replaced their original feature encoding with our own and substituted their scalar output layers with a LogSoftmax layer. Among them, FCNN is the only model that fails to outperform Spark SQL even after training on 2400 queries. Specifically, LSTM exhibits the lowest optimization overhead, making it more efficient during both training and inference. However, TreeCNN demonstrates better convergence behavior and achieves higher overall performance after training, albeit with a slightly increased overhead.
6.5.3. Different Learning Strategy
@@
To evaluate the effectiveness of our learning strategy design, we conducted an ablation study using two alternative strategies (Figure 9(c)). In the first strategy, we removed the step limit during optimization while retaining the curriculum learning schedule. In the second strategy, we limited the maximum number of steps but removed the curriculum constraint, allowing LQRS to explore the full action space from the beginning.
@@
Experimental results show that the first strategy performs comparably to the default configuration during the early training phase. However, as the curriculum progresses to its second stage, the model tends to take significantly more steps, making it difficult to repeatedly make correct decisions. The second strategy, despite enforcing a strict step limit, ultimately fails to outperform Spark SQL. This is because, in most cases, only a few targeted corrections are sufficient to reach an efficient plan. Without early-stage guidance, the model explores an overly large action space too freely, which hinders convergence.
6.5.4. Different Action Space
@@
LQRS’s planner extension supports a wide range of action types. We conducted ablation studies to identify which subsets of the action space contribute most to learning efficiency and convergence. As shown in the results, enabling the broadcast action often leads the model to generate plans that broadcast excessively large tables during training, resulting in significant performance fluctuations and instability. When the lead action is disabled, LQRS loses its ability to explicitly manipulate join ordering, thereby reducing its optimization power and leading to suboptimal plans for many complex queries. Finally, disabling the init action prevents LQRS from jumping to an entirely different planning path starting from a CBO-generated plan. This restriction increases the difficulty for LQRS to escape local optima and discover better execution plans.
6.6. Overhead Study
Table 4. Optimization cost on all benchmarks.
Time (s) JOB ExtJOB STACK TPC-H
Lero 30098.1 201.3 1110.6 379.5
AutoSteer 829.2 176.7 634.0 442.1
SSA 81.5 6.2 41.0 27.7
LQRS 33.3 7.1 37.3 19.6
As shown in Table 4, LQRS achieves the lowest optimization overhead among most evaluated methods, by making re-optimization decisions during query execution based on accurate runtime information. Its design explicitly supports a no-op action, allowing the optimizer to skip unnecessary plan modifications when the expected benefit is limited, which effectively suppresses redundant re-optimizations and keeps the optimization process lightweight.
In comparison, SSA incurs higher overhead because it continuously modifies the plan whenever the current leading leaf node is not estimated as the lowest-cost node, resulting in unnecessary plan modifications during execution. Learned query optimizers such as Lero and AutoSteer suffer from even greater costs due to explicit enumeration and evaluation of a large plan search space before execution. Lero perturbs cardinality estimates at multiple levels to generate numerous alternative plans, while AutoSteer explores different combinations of optimization rules, both requiring repeated cost estimation over multiple candidates.
7. Conclusion
We propose LQRS, a learned query re-optimizer integrated into Spark SQL. LQRS serves as a prototype system that demonstrates how execution-time feedback and intervention can be incorporated into a modern query engine via a planner extension. By continuously observing partial execution results and applying query stage-level feedback with runtime-aware rewards, LQRS incrementally refines the remaining execution plan during query execution. Extensive experiments on Spark SQL demonstrate that LQRS reduces end-to-end execution time by up to 90% compared to other learned query optimizers and query re-optimization methods.